I have been working on these questions and cannot get the answer. Could you kindly help me.
1. a^2-a= -a ( ) complete the factorization of each binomial
2. t^2 - 9w^2 Factor polunomial completely
Thank you kindly.

Answer by edjones(8007)

1. a^2-a= -a(-a+1)
2.t^2 - 9w^2=(t-3w)(t+3w)
Ed